# SkillWrapper: Generative Predicate Invention for Skill Abstraction

### v2

@@ H.5 Language Model Prompts
In this section, we provide the prompts used for the core components of SkillWrapper (specifically skill sequence proposal, predicate invention, and predicate evaluation) as well as the ViLA (hu2023lookleapunveilingpower) baseline. For predicate evaluation (Appendix H.8), we empirically observed that it is more accurate when evaluation is done in batches, where the truth values of multiple predicates are evaluated at once rather than one at a time. In addition, when asking for a fixed and structured output, the accuracy is significantly lower than a free-form output. Therefore, we adopt a two-stage evaluation process: in the first stage, the foundation model generates a response in any format, and in the second stage, it provides a summary of the output from the previous step.
@@ Appendix I Related Works
There has been a long track of works focusing on building hierarchies that abstract away high-dimensional details with low-dimensional abstractions for planning (konidaris2009skill; Konidaris2018FromST; shah2024reals), and those applied to robotics are usually connected to task and motion planning (TAMP) (shah2020anytime; garrett2021integrated). These approaches, however, are incapable of handling high-dimensional sensory-motor signals (such as images) as input. Research on action model learning (xi2024neuro; juba2021safe) learn symbolic action models for input skills. However, unlike our method, these approaches require symbols to be provided as input. Similar to our system’s integration of self-play and focus on uncovering skill conditions, verma2022discovering focus on assessing capabilities of black-box agents for grid world-like tasks while assuming that the agent is an oracle. A tangential research effort on chaining various skills in novel environments involves training extra models (yokoyama2023ascadaptiveskillcoordination) and STRIPS task planner with action primitives (gu2022multiskillmobilemanipulationobject; szot2022habitat20traininghome).
@@
Predicates provide a convenient way to abstract away low-level details of the environment and build efficient and compact representations. Prior to foundation models, previous attempts to build classifiers for predicates from raw image inputs originated from the neuro-symbolic domain (johnson2016clevrdiagnosticdatasetcompositional; mao2019neurosymbolicconceptlearnerinterpreting), and their initial application for robotics took a similar supervised learning approach with labeled demonstrations (migimatsu2022grounding) or generated tasks (Lamanna2023PlanningFL). After the emergence of foundation models, recent works guide skill learning with predicates generated by LLM or together with human interaction. li2024league invents symbolic skills for reward functions used for RL training but cannot generalize to skills learned through latent objectives, which is more commonly seen in imitation learning. Li2023EmbodiedAL and han2024interpret leverage human experts to provide feedback to the LLM to help it improve the learned predicates and skills.
@@
The approach of automatically proposing tasks has been studied for active learning and curriculum learning in grid worlds and games (wang2019paired; jiang2021prioritizedlevelreplay) to robotic domains (fang2021adaptiveproceduraltaskgeneration; Fang2022ActiveTR). Lamanna2023PlanningFL generates tasks in PDDL as training sets to learn classifiers for object properties in predicates format, while they assume the action operators are given. With the commonsense reasoning ability of foundation models, recent works have applied the idea of automatic task proposing and self-playing for exploration (google2024pivot; exploreeqa2024), data collection (wang2023robogen; yang2023octopus; ahn2024autort), boosting skills learning (Ha2023ScalingUA; wang2023voyageropenendedembodiedagent), and scene understanding (jiang2024roboexp). These works indicate a promising direction for generating robotic data and scaling up. Following the idea, we equipped our system with a task-proposing module for generating skill sequences specific to skills and predicates, which serves the idea of both data collection and exploration.
@@
There has been a track of work on leveraging large language models (LLMs) for embodied decision-making (huang2022languagemodelszeroshotplanners; raman2024cape) and reasoning (huang2022inner), while vision-language models (VLMs) are often considered to have limited embodied reasoning ability due to their pre-training corpora that focus primarily on language generation (valmeekam2023llm). Common ways of addressing this issue include fine-tuning on datasets from a specific domains (hong20233dllm; mu2023embodiedgptvisionlanguagepretrainingembodied; chen2024spatialvlm) or knowledge distillation (sumers2023distillinginternetscalevisionlanguagemodels; yang2023octopus). Meanwhile, many works manage to leverage preexisting models without further training from direct visual observation (liu2024moka; google2024pivot) to complete robotic tasks (jiang2024roboexp). In these works, the embodied reasoning ability of the foundation models serves as the central part of the systems. However, most benchmarking works evaluate the embodied reasoning ability of the models in a question-answering fashion (sermanet2023robovqamultimodallonghorizonreasoning; OpenEQA2023; cheng2024egothink; chen2025visionlanguage), where it remains unclear whether they are capable of solving robotic tasks.
@@
Our work incorporates language models as part of SkillWrapper, particularly for the three important components of our system discussed in Section 3. We utilize OpenAI’s GPT-5 (openai_gpt5) as our foundation model of choice. We acknowledge that all the content of the manuscript has been generated by the authors. However, we have used LLMs for basic editing, polishing, and grammar checking.

### v3

@@ H.5 Language Model Prompts
In this section, we provide the prompts used for the core components of SkillWrapper (specifically skill sequence proposal, predicate invention, and predicate evaluation) as well as the ViLA (Hu et al., 2023) baseline. For predicate evaluation (Appendix H.8), we empirically observed that it is more accurate when evaluation is done in batches, where the truth values of multiple predicates are evaluated at once rather than one at a time. In addition, when asking for a fixed and structured output, the accuracy is significantly lower than a free-form output. Therefore, we adopt a two-stage evaluation process: in the first stage, the foundation model generates a response in any format, and in the second stage, it provides a summary of the output from the previous step.
@@ Appendix I Related Works
There has been a long track of works focusing on building hierarchies that abstract away high-dimensional details with low-dimensional abstractions for planning (Konidaris & Barto, 2009; Konidaris et al., 2018; Shah et al., 2024), and those applied to robotics are usually connected to task and motion planning (TAMP) (Shah et al., 2020; Garrett et al., 2021). These approaches, however, are incapable of handling high-dimensional sensory-motor signals (such as images) as input. Research on action model learning (Xi et al., 2024; Juba et al., 2021) learn symbolic action models for input skills. However, unlike our method, these approaches require symbols to be provided as input. Similar to our system’s integration of self-play and focus on uncovering skill conditions, Verma et al. (2022) focus on assessing capabilities of black-box agents for grid world-like tasks while assuming that the agent is an oracle. A tangential research effort on chaining various skills in novel environments involves training extra models (Yokoyama et al., 2024) and STRIPS task planner with action primitives (Gu et al., 2022; Szot et al., 2021).
@@
Predicates provide a convenient way to abstract away low-level details of the environment and build efficient and compact representations. Prior to foundation models, previous attempts to build classifiers for predicates from raw image inputs originated from the neuro-symbolic domain (Johnson et al., 2017; Mao et al., 2019), and their initial application for robotics took a similar supervised learning approach with labeled demonstrations (Migimatsu & Bohg, 2022) or generated tasks (Lamanna et al., 2023). After the emergence of foundation models, recent works guide skill learning with predicates generated by LLM or together with human interaction. Li et al. (2024) invents symbolic skills for reward functions used for RL training but cannot generalize to skills learned through latent objectives, which is more commonly seen in imitation learning. Li & Silver (2023) and Han et al. (2024) leverage human experts to provide feedback to the LLM to help it improve the learned predicates and skills.
@@
The approach of automatically proposing tasks has been studied for active learning and curriculum learning in grid worlds and games (Wang et al., 2019; Jiang et al., 2021) to robotic domains (Fang et al., 2021; 2022). Lamanna et al. (2023) generates tasks in PDDL as training sets to learn classifiers for object properties in predicates format, while they assume the action operators are given. With the commonsense reasoning ability of foundation models, recent works have applied the idea of automatic task proposing and self-playing for exploration (Nasiriany et al., 2024; Ren et al., 2024), data collection (Wang et al., 2024c; Yang et al., 2024; Ahn et al., 2024), boosting skills learning (Ha et al., 2023; Wang et al., 2024a), and scene understanding (Jiang et al., 2025). These works indicate a promising direction for generating robotic data and scaling up. Following the idea, we equipped our system with a task-proposing module for generating skill sequences specific to skills and predicates, which serves the idea of both data collection and exploration.
@@
There has been a track of work on leveraging large language models (LLMs) for embodied decision-making (Huang et al., 2022; Raman et al., 2024) and reasoning (Huang et al., 2023), while vision-language models (VLMs) are often considered to have limited embodied reasoning ability due to their pre-training corpora that focus primarily on language generation (Valmeekam et al., 2023). Common ways of addressing this issue include fine-tuning on datasets from a specific domains (Hong et al., 2023; Mu et al., 2024; Chen et al., 2024) or knowledge distillation (Sumers et al., 2023; Yang et al., 2024). Meanwhile, many works manage to leverage preexisting models without further training from direct visual observation (Fang et al., 2024; Nasiriany et al., 2024) to complete robotic tasks (Jiang et al., 2025). In these works, the embodied reasoning ability of the foundation models serves as the central part of the systems. However, most benchmarking works evaluate the embodied reasoning ability of the models in a question-answering fashion (Sermanet et al., 2024; Majumdar et al., 2024; Cheng et al., 2024; Chen et al., 2025), where it remains unclear whether they are capable of solving robotic tasks.
@@
Our work incorporates language models as part of SkillWrapper, particularly for the three important components of our system discussed in Section 3. We utilize OpenAI’s GPT-5 (OpenAI, 2025) as our foundation model of choice. We acknowledge that all the content of the manuscript has been generated by the authors. However, we have used LLMs for basic editing, polishing, and grammar checking.